# Visual Alignment of Medical Vision-Language Models for Grounded Radiology Report Generation

### v1

@@ Abstract
Radiology Report Generation (RRG) is a critical step toward automating healthcare workflows, facilitating accurate patient assessments, and reducing the workload of medical professionals. Despite recent progress in Large Medical Vision-Language Models (Med-VLMs), generating radiology reports that are both visually grounded and clinically accurate remains a significant challenge. Existing approaches often rely on large labeled corpora for pre-training, costly task-specific preference data, or retrieval-based methods. However, these strategies do not adequately mitigate hallucinations arising from poor cross-modal alignment between visual and linguistic representations. To address these limitations, we propose VALOR: Visual Alignment of Medical Vision-Language Models for GrOunded Radiology Report Generation. Our method introduces a reinforcement learning–based post-alignment framework utilizing Group-Relative Proximal Optimization (GRPO). The training proceeds in two stages: (1) improving the Med-VLM with textual rewards to encourage clinically precise terminology, and (2) aligning the vision projection module of the textually grounded model with disease findings, thereby guiding attention toward image regions most relevant to the diagnostic task. Extensive experiments on multiple benchmarks demonstrate that VALOR substantially improves factual accuracy and visual grounding, achieving significant performance gains over state-of-the-art report generation methods.
Figure 1: Overview: We introduce VALOR, a multimodal alignment framework that addresses the visual hallucinations of existing Med-VLMs [li2023llava]. Left: the generated response from vanilla Med-VLM where the misalignment with the input image is marked in red and generic non-medical text in cyan. Existing approaches resort to the curation of preference data per-task or on retrieval of reports which is inherently time-consuming and does not address the issue of multimodal alignment. Right: Our proposed VALOR addresses this issue with a novel visual reasoning based pipeline which utilizes multi-modal rewards for optimization and pushes the model to generate more clinical informed and grounded reports aligned with the original image which are marked in green.
[…]
@@ 1 Introduction
Radiology reports constitute a fundamental component of assistive diagnosis, serving as a structured medium through which complex visual findings in X-ray images are translated into clinically actionable insights. By standardizing the description of anatomical regions and pathological observations, these reports not only enable more efficient and consistent patient assessment but also alleviate the cognitive and administrative burdens on healthcare professionals.
@@
Figure 2: Comparison with SOTA methods. VALOR performs significantly better in both generation [papineni2002bleu, lin2004rouge, denkowski2011meteor] and clinical metrics [smit2020chexbert, jain2021radgraph] compared to preference-data and retrieval-based methods. Further, they don’t account for visual grounding w.r.t the input image as shown in the table.
Preparing chest X-ray reports requires deep clinical and anatomical expertise to visually identify disease-relevant regions and accurately describe them. This burden has motivated the development of automated chest X-ray report generation techniques, largely driven by the rise of Large Medical Vision–Language Models (Med-VLMs) [li2023llava, zhou2023skingpt, zhou2023path, sellergren2025medgemma, nath2025vila]. However, improving cross-modal alignment [amirloo2024understanding, cao2024madtp, zhou2024aligning]–that is, the ability to explicitly associate radiological findings with corresponding X-ray regions, and thereby ground generated reports–is critical to ensure that these reports contain image-relevant disease findings. Failure to do so results in “Visual Hallucinations” stemming from the inherent tendency of these models to generate information from their reliance on a text-heavy pre-training corpus [leng2024mitigating, Huang_2024_CVPR, liu2025gemex] and using retrieval-augmented pipelines that may incorporate disease statements from retrieved reports even when unsupported by the image [xia2024rule, hou-etal-2025-radar] bypassing actual multimodal reasoning [chen2024detecting] needed for explainability of report generation systems, as shown in Figure Visual Alignment of Medical Vision-Language Models for Grounded Radiology Report Generation.
To mitigate these challenges, existing efforts can be broadly divided into two categories: 1) Construction of high-quality medical preference datasets: either by leveraging expert preferences [chen2024huatuogpt, cui2024biomedical] or by constructing additional data generation and training pipelines requiring significant time and extra computational resources [zhu2024mmedpo]. While promising, these methods remain overly sensitive to the quality of preference data, often generated using simplistic techniques such as random masking or additive noise, which may result in loss of information and, most importantly, fail to explicitly enforce that generated reports are grounded in the input X-ray. 2) Domain Knowledge Retrieval: by either training domain-specific retrievers or curating retrieval knowledge bases [xia2024rule, sun2025fact, hou-etal-2025-radar]. However, the scarcity of such external multimodal medical knowledge, the inherent semantic gap between retrieved reports and medical images, and the limited scope of the retrieval set all lead to ineffective visual grounding as reflected in NLG and Clinical Metrics [papineni2002bleu, lin2004rouge, denkowski2011meteor, smit2020chexbert] (Figure 2). These prompts us to ask the question: Can we design a framework that generates clinically informative reports which are both factually accurate and visually grounded in the image?
To address these challenges, we propose VALOR: Visual Alignment of Medical Vision-Language Models for GrOunded Radiology Report Generation, a post-training alignment framework that improves the medical reasoning of Med-VLMs with both textual and visual rewards to reduce hallucinations and strengthen image–report consistency. Our main contributions are threefold:
  • Multi-Stage Alignment. We introduce a novel multi-stage reinforcement learning-based pipeline, where the base Med-VLM [li2023llava] is aligned in the textual space and then refined for visual alignment. In the first stage, we craft textual rewards with natural language metrics BLEU [papineni2002bleu], ROUGE [lin2004rouge], BERTScore [zhang2019bertscore] for semantic completeness, and clinical metrics CheXbert [smit2020chexbert] and RadGraph [jain2021radgraph] for clinical guidance, with format consistency rewards. In the second stage, we utilize multi-label features to compute image–text similarity scores, explicitly steering the model towards disease-relevant regions.
  • Hybrid Reward Formulation for Reasoning. To achieve multimodal medical reasoning, our novel multi-stage reward regimen blends commonly used NLG and Clinical metrics as rewards to supervise the model with enriched medical terminology. We then use normalized image-text similarity scores between domain features from a disease classifier [irvin2019chexpert] and generated candidate reports to compute rank-normalized advantages, assigning more weightage to reports that are aligned to the image, thus enabling effective visual-text integration.
  • Extensive evaluation on Radiology Report Generation. We extensively evaluate VALOR on the IU-XRay [demner2015preparing] and MIMIC-CXR [johnson2020mimic] datasets across diverse scenarios: generation quality, clinical accuracy and verify that our method achieves effective visual grounding qualitatively. VALOR consistently outperforms preference-based, retrieval-based and SOTA proprietary LLMs producing more faithful reports whose attention maps concentrate on clinically meaningful regions, demonstrating improved factual accuracy and visual alignment.
@@ 2 Related Works
Radiology Report Generation (RRG). Radiology report generation involves composing descriptive sentences that capture the contents of a medical image, organized into a ‘Findings’ section for detailed observations and an Impression section for clinical interpretation. While standard image captioning tasks involve answering in brief, closed-set settings, RRG typically involves generating much larger quantities of text in a precise manner that is clinically informed and aligned with the image. Various works have been proposed along these lines [huang2023kiut, jin2024promptmrg, li2023dynamic].
@@
Existing approaches for RRG can be broadly categorized into three groups. The first focuses on directly improving the encoder-decoder architecture to generate more consistent and medically enriched reports. Earlier approaches used LSTM [hochreiter1997long] based networks with hierarchical structures to manage the descriptiveness of radiology reports [tanida2023interactive] or dynamic graphs with a contrastive learning paradigm [li2023dynamic]. The second group of approaches focuses on closest report retrieval using retrieval-augmented generation (RAG) techniques, where a retriever is trained based on medical informativeness and targeted retrieval using domain knowledge [li2024rag, dong2025understand, xia2024mmed, chu2025reducing]. RADAR [hou-etal-2025-radar] employs the combination of retrieved reports and the overlap with the potential diseases which may be there in the image. However, these methods employ terminology that is misaligned with the image, limiting their reliability in clinical practice.
@@
Reinforcement Learning in LLMs/VLMs. Despite the rise of Supervised Fine-Tuning (SFT)-based techniques, Reinforcement Learning (RL)-based approaches have also gained popularity owing to their robustness, scalability, ability to align with human preferences, and ability to mitigate hallucinations [yu2025rlaif, zhou2024aligning, wang2025instruction]. MMedPO [zhu2024mmedpo] utilizes a Direct Preference Optimization (DPO)-based approach based on curating preference data to mitigate hallucinations in Med-VLMs. However, it requires the curation of preference data, which is time-consuming and expensive, and does not generalize well outside of the curated preference data. Our work focuses on addressing this limitation by applying GRPO [shao2024deepseekmathpushinglimitsmathematical]-based RL using verifiable rewards with a focus on unifying both modalities, which enhances the medical reasoning capabilities of the model.
@@
Medical Reasoning using Vision Language Models. With the widespread success of CLIP [radford2021learning], several other VLMs have arisen in the medical space such as Med-Flamingo [moor2023med], BioMedGPT [zhang2024generalist] and HuatuoGPT-Vision [chen2024huatuogpt]which learn the cross-modal information across a large scale of pre-training data across various vision and language modalities such as CheXzero [tiu2022expert], Med-CLIP [wang2022medclip] and BioMedCLIP [zhang2023biomedclip] which focus on enhancing the understanding of raw reports with images whereas BioVIL [boecking2022making], BioVIL-T [bannur2023learning] enhance the encoding of the textual descriptions. More recent work, such as MedKLIP [wu2023medklip] leverages the spatial relationships from RadGraph [jain2021radgraph] to incorporate more fine-grained medical information into the pre-training regimen. DART [park2025dart] proposed a self-correcting report generation mechanism that focused on improving both the encoding and decoding processes. Methods such as MMedAgent [li2024mmedagent] and VILA-M3 [nath2025vila] utilize expert knowledge (requiring annotation) and tool usage (prone to model noise) as additional proxies of information for alignment to the medical domain. Thus, these successes can be largely attributed to the availability of a large corpus of labeled pretrained data and largely ignore the alignment of human preferences with the provided image. It is therefore important to move away from external dependencies such as tools, preference data or retrieval datasets and adjust the inherent reasoning focus of the model towards improved clinical accuracy and groundedness to the input X-Ray. Our novel multi-stage RL-based reasoning pipeline aims to address this open challenge of improving the visual abilities of Med-VLMs without using additional external knowledge.
@@ 3 Preliminaries
Figure 3: Workflow of VALOR. Left: In Stage-1, VALOR is trained with verifiable textual rewards $R_{ver}$ with periodic clinical guidance $R_{clin}$ every $k_{clin}$ policy steps to familiarize the base Med-VLM $\mathcal{M(\cdot)}$ [li2023llava] with radiology reasoning patterns. Right: In Stage-2, VALOR is further optimized with rewards aligned using image-text similarity scores computed by using embeddings obtained utilizing the multi-label domain expert model $\mathcal{G}$ [irvin2019chexpert], shifting the model’s reasoning focus from textual knowledge, to grounded, region-wise disease reasoning capabilities. We also utilize format rewards to maintain structure of generated reports using <report> tags.
Problem Formulation. Given a 2D chest X-ray image $\mathcal{I}\in\mathbb{R}^{H\times W\times 3}$, a textual prompt $P$, and a dataset $\mathcal{D}=\{(I,\mathcal{R}_{\text{gt}})\}$ of images and ground truth reports, the goal of radiology report generation is to produce a report $\mathcal{R}$ that is clinically accurate and aligned with the input image $I$. We model this with a base Med-VLM $\mathcal{M}_{\theta}$, which induces a conditional policy $\pi_{\theta}(R\mid\mathcal{I})$ that generates $\mathcal{R}$ in an autoregressive manner. In this work, we adapt $\mathcal{M}_{\theta}$ in two stages: first, we refine the text generation policy using verifiable and clinical textual rewards; second, we further align the model with image-conditioned visual rewards derived from a frozen domain expert, encouraging reports whose disease findings are grounded in the underlying X-ray $\mathcal{I}$.
RRG Objective. Radiology Report Generation refers to the generation of structured, coherent, and medically informed text from a medical image. In this work, we follow existing works and process the existing report in two parts: the FINDINGS and IMPRESSION sections, where the former contains the broad rationale behind the generated report and the latter contains the potential diseases present in the patient’s medical image. Formally, this process can be represented as follows: given the medical image $\mathcal{I}$, the Med-VLM $\mathcal{M}(\cdot)$ is tasked with interpreting the image, reasoning about it, and generating a descriptive report $\mathcal{R}=\{r_{1},r_{2},\ldots,r_{T}\}$ where $T$ represents the total length of the report and $r_{t}\in\mathbb{V}$ represents the tokens in the vocabulary. The report $\mathcal{R}$ is auto-regressively generated as,
$p(\mathcal{R}\mid I)=\prod_{t=1}^{T}p(r_{t}\mid r_{1},\ldots,r_{t-1},I).$ (1)

### v2

@@ Abstract
Radiology Report Generation (RRG) is a critical step toward automating healthcare workflows, facilitating accurate patient assessments, and reducing the workload of medical professionals. Despite recent progress in Large Medical Vision-Language Models (Med-VLMs), generating radiology reports that are both visually grounded and clinically accurate remains a significant challenge. Existing approaches often rely on large labeled corpora for pre-training, costly task-specific preference data, or retrieval-based knowledge. However, these strategies do not adequately mitigate hallucinations arising from poor cross-modal alignment between visual and linguistic representations. To address these limitations, we propose VALOR: Visual Alignment of Medical Vision-Language Models for GrOunded Radiology Report Generation, which tackles visual hallucinations through two complementary reasoning stages: (1) Clinically Informed Textual Reasoning guides the model with verifiable natural language and clinical metric rewards to produce semantically complete reports with precise medical terminology. (2) Self-Supervised Visual Reasoning leverages a frozen domain expert to compute image-text similarity scores between the input chest X-ray and generated candidates, converting these into rank-normalized advantages that explicitly steer the policy toward visually grounded outputs, requiring no preference pairs, retrieval databases, or additional annotations. Extensive experiments on multiple benchmarks demonstrate that VALOR substantially improves generation quality, as well as clinical accuracy which are visually grounded, achieving significant performance gains over state-of-the-art medical report generation benchmarks.
Figure 1: Overview: We introduce VALOR, a multimodal alignment framework that addresses the visual hallucinations of existing Med-VLMs [li2023llava]. Left: the generated response from vanilla Med-VLM where the misalignment with the input image is marked in red and generic non-medical text in cyan. Existing approaches resort to curation of preference data per-task or on retrieval of reports which is inherently expensive and does not address the issue of multimodal alignment. Right: Our method VALOR, addresses this issue with a novel visual reasoning based pipeline which utilizes multi-modal rewards for optimization and pushes the model to generate more clinically informed and grounded reports aligned with the original image which are marked in green.
@@ 1 Introduction
Automated chest X-ray report generation has emerged as a promising avenue to reduce the burden of medical professionals for manual radiology reporting. This has emerged with the rise of Large Medical Vision-Language Models (Med-VLMs) [li2023llava, zhou2023skingpt, zhou2023path, sellergren2025medgemma, nath2025vila] becoming the dominant paradigm, leveraging large-scale multimodal pretraining and external clinical knowledge to generate reports directly from images. However, these models frequently produce reports that are not grounded in the input image, generating clinically plausible but unsupported findings, a phenomenon known as “Visual Hallucination” (Fig. Visual Alignment of Medical Vision-Language Models for Grounded Radiology Report Generation). This arises primarily because Med-VLMs tend to rely on learned patterns arising from text-heavy pretraining corpora rather than performing genuine cross-modal reasoning [leng2024mitigating, Huang_2024_CVPR, liu2025gemex, amirloo2024understanding, cao2024madtp, zhou2024aligning]. The problem is further compounded in retrieval-augmented pipelines, where clinical statements from retrieved reports can be incorporated even when they are inconsistent with the visual evidence [xia2024rule, hou-etal-2025-radar, chen2024detecting]. This prompts a fundamental question: Can we design a framework that generates clinically informative reports which are both factually accurate and visually grounded in the image?
Existing efforts to mitigate visual hallucinations face two fundamental limitations: (1) Data Reliance: Methods that rely on preference datasets [chen2024huatuogpt, cui2024biomedical, ouyang2022training, zhu2024mmedpo] are inherently bounded by the quality of their preference pairs, which are typically constructed using simplistic techniques devoid of clinical context. More critically, the preference optimization usually happens entirely in the textual space: the model learns to prefer one report over another without any mechanism to verify whether either report is actually supported by the input image and does not account for the inherent gap between reports and provided images (2) Visual Misalignment: Retrieval-augmented approaches [jin2021disease, xia2024rule, sun2025fact, hou-etal-2025-radar] inject external clinical context, but introduces a new source of hallucination, since retrieved reports describe other patients’ images and may contain findings entirely absent from the current X-ray (Fig. Visual Alignment of Medical Vision-Language Models for Grounded Radiology Report Generation). Further, this requires training retrieval models on domain specific data which adds computational overhead. The semantic gap between retrieved text and the visual evidence is fundamental, not merely a matter of retrieval quality. As a result, neither paradigm enforces what is most needed: direct alignment between the generated text and the visual content of the image, without relying on external preference data, retrieval databases, or additional annotations (Fig. 2).
@@
Figure 2: Comparison with SOTA methods. VALOR performs significantly better in generation quality [papineni2002bleu, lin2004rouge, denkowski2011meteor] and clinical accuracy [irvin2019chexpert, ostmeier2024green, zhao2024ratescore] without using any preference-data or additional retrieval-based methods. Further, these works do not perform disease-specific visual grounding w.r.t. the given image as shown in the table.
To address these challenges, we propose VALOR: Visual Alignment of Medical Vision-Language Models for GrOunded Radiology Report Generation. VALOR enhances the medical reasoning of Med-VLMs through a two-stage multimodal optimization process. In the first stage, Clinically Informed Textual Reasoning, the model is guided by verifiable natural language generation metrics with explicit clinical guidance as rewards to produce semantically coherent reports with precise clinical terminology. In the second stage, Self-Supervised Visual Reasoning, we leverage a frozen multi-label domain expert model to compute image-text similarity scores in-between the input X-ray and each generated candidate reports per rollout, converting these into rank-normalized advantages that explicitly steer the policy toward visually grounded outputs. In summary, our main contributions are as follows:
  1. 1.
    We propose VALOR, a novel framework that generates visually grounded and clinically informative radiology reports through two complementary reasoning stages, reducing visual hallucinations and strengthening image-report consistency without relying on any external curated preference datasets or domain specific retrieval-based methods.
  2. 2.
    We introduce clinically-informed textual rewards for precise terminology in the generated report and rank-normalized advantages, which utilizes image-text similarity scores obtained from a frozen domain expert for visual guidance, allowing the policy to distinguish which candidate reports are closer to the input X-ray. This enables visual grounding without requiring preference pairs, retrieval databases, or additional annotations, using only the agreement between the generated report and disease-relevant features for grounded report generation.
  3. 3.
    We conduct extensive evaluations of VALOR on IU-XRay [demner2015preparing] and MIMIC-CXR [johnson2020mimic], reporting both generation quality and clinical accuracy metrics, which are rarely assessed jointly in prior work. We also provide a qualitative visual grounding analysis through attention visualizations. As shown in Fig. 2, VALOR consistently outperforms both baseline and state-of-the-art preference-alignment methods by approximately 20% in generation quality and 25% in clinical accuracy. Furthermore, in a zero-shot setting, VALOR surpasses medical and proprietary Med-VLMs by approximately 17%, generating more coherent and visually grounded reports that focus on clinically meaningful regions with improved factual accuracy.
@@ 2 Related Works
Radiology Report Generation (RRG). Radiology report generation involves composing descriptive sentences that capture the contents of a medical image, organized into a “Findings” section for detailed observations and an “Impression” section for fine-grained interpretation that is clinically informed and aligned with the image. Various works have been proposed along these lines [huang2023kiut, jin2024promptmrg, li2023dynamic]. Existing approaches for RRG can be broadly categorized into three groups. The first focuses on directly improving the encoder-decoder architecture to generate more consistent and medically enriched reports. Earlier approaches used LSTM [hochreiter1997long] based networks with hierarchical structures to manage the descriptiveness of radiology reports [tanida2023interactive], dynamic graphs with a contrastive learning paradigm [li2023dynamic] or supervised multimodal pre-training [tanida2023interactive, wang2023metransformer, park2025dart]. The second group of approaches focuses on closest report retrieval using retrieval-augmented generation (RAG) techniques, where a retriever is trained based on medical informativeness and targeted retrieval using domain knowledge [li2024rag, dong2025understand, xia2024mmed, chu2025reducing]. RADAR [hou-etal-2025-radar] employs the combination of retrieved reports and the overlap with the potential diseases which may be there in the image. However, these methods either employ paired multimodal data, require additional retrieved information during training or inference and generate terminology that is misaligned with the image, limiting their reliability in clinical practice.
@@
Reinforcement Learning in LLMs/VLMs. Despite the rise of Supervised Fine-Tuning (SFT)-based techniques, Reinforcement Learning (RL)-based approaches have also gained popularity owing to their robustness, scalability, ability to align with human preferences, and ability to mitigate hallucinations [yu2025rlaif, zhou2024aligning, wang2025instruction]. MMedPO [zhu2024mmedpo] utilizes a Direct Preference Optimization (DPO)-based approach based on curating preference data to mitigate hallucinations in Med-VLMs. However, it requires the curation of preference data, which is time-consuming and expensive, and does not generalize well outside of the curated preference data. Our work focuses on addressing this limitation by proposing a novel two-stage post-training recipe that focus on unifying both modalities, and progressively enhances the visual and clinical reasoning capabilities of the model.
@@
Medical Reasoning using Vision Language Models. With the widespread success of CLIP [radford2021learning], several other VLMs have arisen in the medical space such as Med-Flamingo [moor2023med], BioMedGPT [zhang2024generalist] and HuatuoGPT-Vision [chen2024huatuogpt]which learn the cross-modal information across a large scale of pre-training data across various vision and language modalities such as CheXzero [tiu2022expert], Med-CLIP [wang2022medclip] and BioMedCLIP [zhang2023biomedclip] which focus on enhancing the understanding of raw reports with images whereas BioVIL [boecking2022making], BioVIL-T [bannur2023learning] enhance the encoding of the textual descriptions. More recent work, such as MedKLIP [wu2023medklip] leverages the spatial relationships from RadGraph [jain2021radgraph] to incorporate more fine-grained medical information into the pre-training regimen. Methods such as MMedAgent [li2024mmedagent] and VILA-M3 [nath2025vila] utilize expert knowledge (requiring annotation) and tool usage (prone to model noise) as additional proxies of information for alignment to the medical domain. Thus, these successes can be largely attributed to the availability of a large corpus of labeled pretrained data and largely ignore the alignment of human preferences with the provided image. It is therefore important to move away from external dependencies such as tools, preference data or retrieval datasets and adjust the inherent reasoning focus of the model towards the provided X-ray. Our novel post-alignment framework aims to address this open challenge of improving the visual reasoning abilities of Med-VLMs without using additional external knowledge in a self-supervised fashion.
@@ 3 Preliminaries
Problem Formulation. Given a 2D chest X-ray image $\mathcal{I}\in\mathbb{R}^{H\times W\times 3}$, a text prompt $\mathcal{P}$, and a dataset $\mathcal{D}=\{(\mathcal{I},\mathcal{R}_{\text{gt}})\}$ of images and ground truth reports, the goal of radiology report generation is to produce a report $\mathcal{R}$ that is clinically accurate and aligned with the input image $\mathcal{I}$. We model this with a base Med-VLM $\mathcal{M}_{\theta}$, which induces a conditional policy over it’s parameters $\pi_{\theta}(\mathcal{R}\mid\mathcal{I})$ to generate $\mathcal{R}$ in an autoregressive manner. In this work, we adapt $\mathcal{M}_{\theta}$ in two stages: first, we refine the text generation policy using verifiable and clinical textual rewards; second, we further align the model with image-conditioned visual rewards derived from a frozen domain expert $\mathcal{G}$, encouraging reports whose disease findings are grounded in the underlying X-ray $\mathcal{I}$.
RRG Objective. We follow existing works and process reports in two parts: the FINDINGS and IMPRESSION sections, where the former contains the broad rationale behind the generated report and the latter contains the potential diseases present in the patient’s medical image. Formally, this process can be represented as follows: given the medical image $\mathcal{I}$, the Med-VLM $\mathcal{M}(\cdot)$ is tasked with interpreting the image, reasoning about it, and generating a descriptive report $\mathcal{R}=\{r_{1},r_{2},\ldots,r_{T}\}$ where $T$ represents the total length of the report and $r_{t}\in\mathbb{V}$ represents the tokens in the vocabulary. The report $\mathcal{R}$ is auto-regressively generated as,
$p(\mathcal{R}\mid\mathcal{I})=\prod_{t=1}^{T}p(r_{t}\mid r_{1},\ldots,r_{t-1},\mathcal{I}).$ (1)
Figure 3: Workflow of VALOR. Left: In Stage-1, VALOR is trained with verifiable textual rewards $R_{ver}$ with periodic clinical guidance $R_{clin}$ every $k_{clin}$ policy steps to familiarize the base Med-VLM $\mathcal{M(\cdot)}$ with radiology reasoning patterns. Right: In Stage-2, VALOR is further optimized with rewards aligned using image-text similarity scores computed by using embeddings obtained utilizing the multi-label domain expert model $\mathcal{G}$, shifting the model’s reasoning focus from textual knowledge, to grounded, region-wise disease reasoning capabilities in a self-supervised fashion. We also utilize format rewards to maintain structure of generated reports using <report> tags.